solve the system
4x+2y=4
6x-8=-2y

Answer by stanbon(75887)

solve the system
4x+2y=4
6x-8=-2y
-------------------------
Rearrange:
4x + 2y = 4
6x + 2y = 8
---------------
Subtract 1st from 2nd and solve for "x":
2x = 4
x = 2
---------------
Substitute to solve for "y":
4*2 + 2y = 4
2y = -4
y = -2
====================
Cheers,
Stan H.